OpenStudy (sophadof):

I need help with a piecewise function(: f(x)=1/4x+5/4 if x x

OpenStudy (johnweldon1993):

So now what are we supposed to do?

OpenStudy (sophadof):

I have no idea...hahah @johnweldon1993

OpenStudy (sophadof):

They give me answers either a,b,c, or d...

OpenStudy (solomonzelman):

I would assume you need to find values of a and b that would make this piece wise function continuous.

OpenStudy (solomonzelman):

But then you have a domain error, and that is because x=a and x=b are excluded from the domain, so you need \(\le\) and \(\ge\) in correct places.

OpenStudy (solomonzelman):

it doesn't matter which part includes a and which part includes b in the domain, but would be nice you post the correct version of the peace wise function.

OpenStudy (solomonzelman):

But, in any case, if I ignore this error for now, but the following 2 statements must be true. at x=a, must be that: 1/4x+5/4 = 2x+3 at x=b, must be that: 2x+3 = -3x+18

OpenStudy (sophadof):

I attatched it..does that help? @SolomonZelman

OpenStudy (solomonzelman):

Yes, I knew it and I was exactly correct, whether that is usual for me or not....

OpenStudy (solomonzelman):

yes, just as I assumed, and in any case, now we can surely proceed to me previous comment: For the function to be continous: at x=a, must be that: 1/4x+5/4 = 2x+3 at x=b, must be that: 2x+3 = -3x+18

OpenStudy (solomonzelman):

so, a/4+5/4=2x+3 --> (a+5)/4=2a+3 --> a+5=8a+12 --> -7=7a --> a=-1 and you solve for b